When I move I have kinetic energy
But my question is
When I move but holding a book in my hand
So does that book also have kinetic energy
If yes then is it half of mine KE which I possess
1
Expert's answer
2019-04-02T10:57:25-0400

When any objections moves, it has the kinetic energy, which can be calculated as:


"E=\frac{mv^2}{2}"

If a person is moving while holding a book, both the person and the book will have kinetic energy.

The fraction of total kinetic energy that the book has is:


"k=\frac{E_{book}}{E_{total}}=\frac{E_{book}}{E_{book}+E_{person}}=\frac{m_{book}}{m_{book}+m_{person}}"

If we assume that the mass of the book is 2 kg,and the mass of the human is 70kg, we get


"k=\frac{2}{2+70}=0.028=2.8\%"

So, the book has only 2.8% of total kinetic energy